# MBA-SLAM: Motion Blur Aware Dense Visual SLAM with Radiance Fields Representation

### v1

@@ 4.3 Evaluation on Standard Sharp Datasets
Tracking: The quantitative tracking results are presented in Table IV. Our method outperforms previous state-of-the-art approaches, including both NeRF-based SLAMs (e.g., NICE-SLAM [10], CoSLAM [11], ESLAM [12], and Point-SLAM [26]) and Gaussian Splatting-based SLAMs (e.g., GS-SLAM [14] and SplaTAM [15]), in terms of both tracking accuracy and running speed (as shown in Table VIII). Notably, our CUDA-implemented tracker operates in real time, with the running speed primarily limited by the back-end mapper. The tracking results for office2 in Table IV are not as strong as those for the other sequences due to the reduced number of features in office2, which hampers our feature-based tracker.
Rendering: We also compare the rendering performance of MBA-SLAM with dense visual SLAMs in Table V. The results demonstrate that MBA-SLAM achieves the best performance, outperforming the second-best implicit dense visual SLAM by 0.6 PSNR in the Our-NeRF version and explicit Gaussian SLAMs by 2.7 PSNR in the Our-GS version.
Reconstruction: The quantitative evaluations of 3D mesh reconstruction against other state-of-the-art dense visual SLAMs are presented in Table VI. MBA-SLAM surpasses all other SLAM systems in Precision, Recall, and F1 metrics on average, achieving comparable Depth L1 performance with SplaTAM [15]. Qualitative mesh visualizations are shown in Fig. 9, where we observe that CoSLAM [11] and ESLAM [12] produce plausible geometry, but lack detail, resulting in over-smoothed mesh outcomes. In contrast, SplaTAM [15] delivers more details, albeit with increased artifacts and protrusions. Compared to these methods, MBA-SLAM provides high-fidelity reconstruction results with more accurate geometric details and fewer artifacts.
@@
Tracking: The quantitative results for tracking presented in Table VII demonstrate that MBA-SLAM achieves the best average camera localization performance, with the exception of our NeRF version (best: 2.00, ours: 2.04) in the TUM RGB-D dataset. Additionally, our Gaussian Splatting version outperforms both SplaTAM [15] and GS-SLAM [14]. The “-” symbol indicates that GS-SLAM is not open-sourced, so we are unable to report the results.
@@
Method Tracking Mapping Mapping System ATE Scene
@@
TABLE IX: Runtime and performance on blur dataset #fr1_desk and #ArchViz-1. ✗ denotes failed running deu to code bugs. Our-NeRF* and Ours-GS* indicate we do not model blur formation process, as other methods.
Method fr1_desk ArchViz-1
@@
NICE-SLAM [10] 32.72$\times$200 49.76$\times$60 0.15 2.96
CoSLAM [11] 3.89$\times$20 15.81$\times$20 6.33 2.75 8.36$\times$50 31.00$\times$40 2.39 5.28 22.97
ESLAM [12] 16.09$\times$200 32.49$\times$60 0.31 2.44 10.32$\times$40 36.24$\times$30 2.42 20.21 21.07
PointSLAM [12] 29.91$\times$200 27.84$\times$150 0.17 2.56 30.99$\times$60 67.22$\times$300 0.20 289.56 15.90
Ours-NeRF* 16.56$\times$1 30.07$\times$60 2.22 2.45 10.73$\times$1 13.82$\times$50 1.99 1.77 23.51
Ours-NeRF 39.64$\times$1 44.74$\times$60 1.49 2.08 35.91$\times$1 36.23$\times$50 0.76 0.98 24.77
SplaTAM [15] 10.59$\times$200 12.06$\times$30 0.47 2.58 18.46$\times$60 22.34$\times$80 0.56 36.88 17.69
Ours-GS* 17.15$\times$1 29.87$\times$100 1.34 2.53 22.14$\times$1 15.29$\times$100 0.90 4.58 22.32
Ours-GS 19.10$\times$1 192.71$\times$160 0.13 2.16 23.91$\times$1 54.95$\times$160 0.17 0.75 28.44
TABLE X: Ablation study on custom Radiance Fields or 3DGS Tracker and our Blur Aware Tracker on sharp dataset #Room0 and blur dataset #ArchViz-1. It demonstrates that our tracker benefits both tracking and mapping process compared to tracking with Radiance Fields or Gaussian Splatting, no matter sharp or blurred input.
Method Use Our Room0 ArchViz-1
Blur-Aware Model ATE Image Image Model ATE Image Image
@@
Ours-NeRF 0.71 25.32 0.875 2.46 23.96 0.873
Ours-NeRF 0.34 28.20 0.916 0.98 24.77 0.905
Ours-GS 0.75 31.98 0.963 1.33 26.19 0.893
Ours-GS 0.25 35.68 0.989 0.75 28.45 0.956
TABLE XI: The effect of the number of interpolated virtual images on dataset #ArchViz-1. It demonstrates that the performance saturates as the number increases.
@@ 4.4 Efficiency Evaluations
Table VIII and Table IX report the runtime of MBA-SLAM and state-of-the-art methods on the standard Replica and blurred ArchViz datasets. We also report the scene representation memory usage on Replica. The results in Table VIII indicate that our NeRF version SLAM system achieves near real-time speed at 22.57 FPS compared to other NeRF-based SLAMs, while our Gaussian Splatting version achieves 1.97 FPS, which is slower than GS-SLAM [14] but faster than SplaTAM [15]. Furthermore, our NeRF and Gaussian Splatting versions achieve the best tracking performance among Radiance Fields based and Gaussian Splatting based SLAMs, respectively. For runtime on the blurred dataset presented in Table IX, we additionally report Our-NeRF* and Our-GS*, which indicate that we do not model the blur formation process in our tracking and mapping. This results in faster FPS due to reduced mapping time, but it leads to a significant decrease in tracking accuracy compared to modeling the blur process (i.e., Our-NeRF and Our-GS). The presence of motion blurred images complicates accurate camera pose estimation, highlighting the robustness of our approach. Without modeling blur, MBA-SLAM achieves comparable or better performance in tracking accuracy and system FPS compared with other methods, except for CoSLAM FPS in the fr1_desk sequence. When modeling the blur process, MBA-SLAM demonstrates improved performance in tracking ATE and rendering (e.g., Fig. 5, Fig. 7 and Fig. 8), albeit with slower system FPS due to the increased mapping time required. Despite this, MBA-SLAM consistently provides the capability for real-time tracking.
@@ 4.5 Ablation Study
Effect of Our Tracker. Table X presents the ablation study of integrating our tracker on the Replica Room0 and ArchViz ArchViz-1 datasets, compared with the custom NeRF-based [12] or 3DGS-based [8] tracker. “Model Blur” refers to modeling the blur formation process in both tracking and mapping. The results demonstrate that our tracker significantly enhances tracking performance, thereby improving mapping and rendering performance on both the sharp Room0 and blurred ArchViz-1 datasets. Furthermore, by integrating our CUDA-implemented tracker, the system can achieve real-time tracking speeds, as shown in Table VIII and Table IX.
Number of Virtual Images. We evaluate the effect of the number of interpolated virtual images (i.e. $n$ in Eq. (2)) within the exposure time. For this experiment, we select two sequences from the synthetic ArchViz dataset: ArchViz-1 and ArchViz-3, representing sequences with low and high levels of motion blur, respectively. The experiments are conducted using our Gaussian Splatting version system with a varying number of interpolated virtual images. The tracking ATE, mapping time, and image quality are reported in Table XI. The results show that as the number of virtual images increases, the tracking and rendering performance (i.e., PSNR, SSIM, LPIPS) tend to saturate, while the mapping time continues to increase. Therefore, after considering the trade-off between system performance and processing speed, we choose to use 13 virtual images for our experiments.
@@ Acknowledgments
This work was supported by National Natural Science Foundation of China (62202389). Special thanks to Wenpu Li for his assistance in debugging the Realsense device and collecting data.

### v2

@@ 4.3 Evaluation on Standard Sharp Datasets
Tracking: The quantitative tracking results are presented in Table IV. Our method outperforms most previous state-of-the-art approaches, including both NeRF-based SLAMs (e.g., NICE-SLAM [10], CoSLAM [11], ESLAM [12], and Point-SLAM [26]) and Gaussian Splatting-based SLAMs (e.g., GS-SLAM [14], SplaTAM [15], Photo-SLAM [17] and MonoGS [16], except RTG-SLAM [51]), in terms of both tracking accuracy and running speed (as shown in Table VIII). Notably, our CUDA-implemented tracker operates in real time, with the running speed primarily limited by the back-end mapper. The tracking results for office2 in Table IV are not as strong as those for the other sequences due to the reduced number of features in office2, which hampers our feature-based tracker.
Rendering: We also compare the rendering performance of MBA-SLAM with dense visual SLAMs in Table V. The results demonstrate that MBA-SLAM achieves the second-best performance after Point-SLAM, outperforming the third-best implicit dense visual SLAM by 0.6 PSNR in the Our-NeRF version and delivering superior or comparable results to other 3DGS-SLAM approaches.
Reconstruction: The quantitative evaluations of 3D mesh reconstruction against other state-of-the-art dense visual SLAMs are presented in Table VI. MBA-SLAM surpasses all other SLAM systems except Point-SLAM in Precision, Recall, and F1 metrics on average, achieving comparable Depth L1 performance with SplaTAM [15] and MonoGS [16]. Qualitative mesh visualizations are shown in Fig. 9, where we observe that CoSLAM [11] and ESLAM [12] produce plausible geometry, but lack detail, resulting in over-smoothed mesh outcomes. In contrast, SplaTAM [15], Photo-SLAM [17] and MonoGS [16] deliver more details, albeit with increased artifacts and protrusions. Compared to these methods, MBA-SLAM provides high-fidelity reconstruction results with more accurate geometric details and fewer artifacts.
@@
Tracking: The quantitative tracking results in Table VII demonstrate that MBA-SLAM achieves average camera localization performance comparable to other state-of-the-art methods. Furthermore, our NeRF version outperforms all other approaches on the ScanNet dataset. The “-” symbol indicates that GS-SLAM is not open-sourced, so we are unable to report the results.
@@
Method Tracking Mapping Keyframe System ATE Scene
@@
RTG-SLAM[51] 60.93 $\times$ 1 4.88 $\times$ 100 400 9.90 0.18 52.26 MB
Photo-SLAM [16] 41.18 $\times$ 1 19.40 $\times$ 100 156 6.61 0.33 59.88 MB
MonoGS [16] 2.54 $\times$ 100 26.56 $\times$ 100 308 1.63 0.42 193.71 MB
@@
TABLE IX: Runtime and performance on blur dataset #pub3 and #ArchViz-1. ✗ denotes failed running due to code bugs. Our-NeRF* and Ours-GS* indicate we do not model blur formation process, as other methods.
Method pub3 ArchViz-1
@@
[…]
TABLE X: Ablation studies on the sharp datasets #Room0 and #Office0 compare custom NeRF and 3DGS-based frame-to-map (f2m) tracking with our proposed frame-to-frame (f2f) tracking.
Method Room0 Office0 Tracking
[…]
COSLAM (f2m) [11] 0.70 27.27 0.910 0.324 0.59 34.14 0.961 0.209 16.64
COSLAM-ours (f2f) 0.32 28.86 0.931 0.272 0.34 35.45 0.969 0.171 29.06
ESLAM (f2m) [12] 0.71 25.32 0.875 0.313 0.57 33.71 0.960 0.184 18.25
ESLAM-ours (f2f) 0.34 28.20 0.916 0.265 0.25 35.27 0.969 0.167 29.00
MonoGS (f2m) [16] 0.42 34.55 0.969 0.071 0.45 40.06 0.991 0.051 3.94
MonoGS-ours (f2f) 0.31 35.07 0.976 0.066 0.36 40.23 0.993 0.049 29.21
TABLE XI: Ablation studies on the blur datasets #ArchViz-1 and #ArchViz-2 compare custom NeRF and 3DGS-based frame-to-map (f2m) tracking with our frame-to-frame (f2f) tracking.
Method ArchViz-1 ArchViz-2 Tracking
[…]
COSLAM (f2m) 5.28 22.97 0.841 0.361 4.67 21.98 0.685 0.570 2.39
COSLAM-blur (f2m) 0.87 25.08 0.904 0.292 1.09 26.03 0.866 0.359 0.61
COSLAM-blur-ours (f2f) 0.61 25.43 0.922 0.265 0.71 26.72 0.893 0.314 29.16
ESLAM (f2m) 20.12 21.07 0.766 0.446 12.61 23.87 0.785 0.475 2.42
ESLAM-blur (f2m) 2.46 23.96 0.873 0.328 3.59 25.23 0.831 0.406 0.57
ESLAM-blur-ours (f2f) 0.98 24.77 0.905 0.312 1.13 26.35 0.875 0.372 27.85
MonoGS (f2m) 1.92 23.18 0.852 0.337 2.96 27.14 0.884 0.329 1.88
MonoGS-blur (f2m) 1.06 27.30 0.924 0.214 0.52 29.92 0.948 0.208 0.41
MonoGS-blur-ours (f2f) 0.68 28.27 0.953 0.148 0.30 30.51 0.960 0.166 41.23
TABLE XII: The effect of the number of interpolated virtual images on dataset #ArchViz-1. It demonstrates that the performance saturates as the number increases.
@@ 4.4 Efficiency Evaluations
Table VIII and Table IX report the runtime of MBA-SLAM and state-of-the-art methods on the standard Replica and blurred ArchViz datasets. We also report the scene representation memory usage on Replica. The results in Table VIII indicate that our NeRF version SLAM system achieves near real-time speed at 22.57 FPS compared to other NeRF-based SLAMs, while our Gaussian Splatting version achieves 1.97 FPS, which is slower than GS-SLAM [14] but faster than SplaTAM [15]. Furthermore, our NeRF and Gaussian Splatting versions achieve the best tracking performance among Radiance Fields based and Gaussian Splatting based SLAMs, respectively. For runtime on the blurred dataset presented in Table IX, we additionally report Our-NeRF* and Our-GS*, which indicate that we do not model the blur formation process in our tracking and mapping. This results in faster FPS due to reduced mapping time, but it leads to a significant decrease in tracking accuracy compared to modeling the blur process (i.e., Our-NeRF and Our-GS). The presence of motion blurred images complicates accurate camera pose estimation, highlighting the robustness of our approach. Without modeling blur, MBA-SLAM achieves comparable or better performance in tracking accuracy and system FPS compared with other methods, except for CoSLAM FPS in the pub3 sequence. When modeling the blur process, MBA-SLAM demonstrates improved performance in tracking ATE and rendering (e.g., Fig. 5, Fig. 7 and Fig. 8), albeit with slower system FPS due to the increased mapping time required. Despite this, MBA-SLAM consistently provides the capability for real-time tracking.
@@ 4.5 Ablation Study
Effect of Our Tracker. Table X and Table XI present the ablation study of integrating our tracker on the Replica Room0, office0 and ArchViz ArchViz-1, ArchViz-2 datasets, compared with the frame-to-frame tracking methods: hash-grid based CoSLAM, tri-plane-based ESLAM and 3DGS-based MonoGS. The results demonstrate that our fully CUDA-implemented frame-to-frame tracker significantly enhances both tracking performance (in terms of ATE and speed) and mapping quality (e.g., PSNR) on both sharp and blurred datasets, compared with NeRF/3DGS-based frame-to-map tracking methods. Explanation of Table X and Table XI: CoSLAM, ESLAM, and MonoGS are the original frame-to-map models. In CoSLAM-ours, ESLAM-ours, and MonoGS-ours, we replace the tracking modules with our frame-to-frame approach while retaining the original mapping. For fair comparison, we add a physical motion blur model to both tracking and mapping, resulting in COSLAM-blur, ESLAM-blur, and MonoGS-blur. Correspondingly, COSLAM-blur-ours, ESLAM-blur-ours, and MonoGS-blur-ours are the blur-aware variants that use our frame-to-frame tracking while preserving the blur-aware mapping components. Furthermore, by integrating our CUDA-implemented tracker, the system can achieve real-time tracking speeds, as shown in Table VIII and Table IX.
Number of Virtual Images. We evaluate the effect of the number of interpolated virtual images (i.e. $n$ in Eq. (2)) within the exposure time. For this experiment, we select two sequences from the synthetic ArchViz dataset: ArchViz-1 and ArchViz-3, representing sequences with low and high levels of motion blur, respectively. The experiments are conducted using our Gaussian Splatting version system with a varying number of interpolated virtual images. The tracking ATE, mapping time, and image quality are reported in Table XII. The results show that as the number of virtual images increases, the tracking and rendering performance (i.e., PSNR, SSIM, LPIPS) tend to saturate, while the mapping time continues to increase. Therefore, after considering the trade-off between system performance and processing speed, we choose to use 13 virtual images for our experiments.
NeRF or 3DGS SLAM. Our NeRF-based SLAM achieves substantially faster training and inference by sampling only a subset of image pixels during both tracking and mapping. This efficiency, however, comes at the cost of reduced rendering fidelity. In contrast, 3DGS-based SLAM employs full-image rendering with Gaussian representations, leading to slower training but higher rendering quality and faster test-time rendering. As shown in Fig. 5, Table II, and Table IX, NeRF-SLAM is preferable for speed-critical applications, while 3DGS-SLAM is more suitable for tasks requiring high-fidelity reconstruction.
@@ Acknowledgments
This work was supported by National Natural Science Foundation of China (62202389).